OpenStudy (anonymous):

What is the value of expression? 6−7 • 64 A. 6−3 B. 6−11 C. 6−28 D. 611

OpenStudy (anonymous):

D. because value just means the answer

OpenStudy (anonymous):

i have 4 more questions care to help me?

OpenStudy (anonymous):

sure

OpenStudy (texaschic101):

is there an exponent in the first problem ?

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

or is it just 6 - 7 * 64 ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

its 7

OpenStudy (texaschic101):

is it 7 or -7 ?

OpenStudy (anonymous):

-7

OpenStudy (texaschic101):

6^-7 * 64 = 1/6^7 * 64 = 1/279936 * 64 = 64/279936 = ....... wait....is the problem 6^-7 * 6^4 ??

OpenStudy (anonymous):

6-^7 the 7 is higher i think the power and also 6-^4

OpenStudy (anonymous):

What is the value of expression? 6−7 • 64

OpenStudy (texaschic101):

x^m * x^n = x^(m + n) 6^-7 * 6^4 = 6^(-7 + 4) = 6^-3

OpenStudy (anonymous):

so its a

OpenStudy (texaschic101):

but if it is : 6^-7 * 6^-4, then it is 6^(-7-4) = 6^-11

OpenStudy (texaschic101):

is the exponent 4 positive or negative

OpenStudy (anonymous):

can u help me with a few more?

OpenStudy (texaschic101):

possibly....not the best at exponents

OpenStudy (anonymous):

that's whjy im here XD

OpenStudy (anonymous):

why A. 7 B. 3 C. −3 D. −7

OpenStudy (texaschic101):

what ??

OpenStudy (anonymous):

What is the missing value? A. 7 B. 3 C. −3 D. −7

OpenStudy (texaschic101):

the missing value in what ?

OpenStudy (anonymous):